OpenStudy (anonymous):

A jar contains 13yellow and 10green marbles. A marble is drawn at random. P(not yellow)

OpenStudy (anonymous):

Ok So there are 23 marbles in all

OpenStudy (anonymous):

so 10/23

OpenStudy (anonymous):

I knew it was jus makin sure thank u

OpenStudy (anonymous):

Probability is a chance game. You have to find your clues and understand them. The jar contains 13 yellow and 10 green. That means that 13+10 = 23 total marbles are in your jar. If a marble is picked randomly, then you have a chance to get either a yellow or a green marble. If there are 13 yellow out of 23 marbles then the chances of getting a yellow is 13/23, similarly there are 10 green out of 23 therefore the chances of getting a green is 10/23. If you get a green marble, then you don't get a yellow marble, and if you get a yellow marble, you don't get a green. The chances of getting a green P(Green) = P(Not yellow) = 10/23